OpenStudy (anonymous):

Find dy/dx by implicit differentiation. x7(x + y) = y2(2x − y)

OpenStudy (anonymous):

\[x^7(x+y)+y^2(2x-4)\]?

OpenStudy (anonymous):

\[\begin{align*}x^7(x+y)&=y^2(2x-y)\\ \frac{d}{dx}\left[x^7(x+y)\right]&=\frac{d}{dx}\left[y^2(2x-y)\right]\\ 7x^6(x+y)+x^7\left(1+\frac{dy}{dx}\right)&=2y\frac{dy}{dx}(2x-y)+y^2\left(2-\frac{dy}{dx}\right)\\ 7x^7+7x^6y+x^7+x^7\frac{dy}{dx}&=\left(4xy-2y^2\right)\frac{dy}{dx}+2y^2-y^2\frac{dy}{dx}\\ 8x^7+7x^6y-2y^2&=\left(4xy-3y^2-x^7\right)\frac{dy}{dx}\\ \frac{dy}{dx}&=\cdots \end{align*}\]